# Refined climatologies of future precipitation over High Mountain Asia using probabilistic ensemble learning

### v1

@@ Abstract
High Mountain Asia holds the largest concentration of frozen water outside the polar regions, serving as a crucial water source for more than 1.9 billion people. In the face of climate change, precipitation represents the largest source of uncertainty for hydrological modelling in this area. Future precipitation predictions remain challenging due to complex orography, lack of in situ hydrological observations, and limitations in climate model resolution and parametrisation for this region. To address the uncertainty posed by these challenges, climate models are often aggregated into multi-model ensembles. While multi-model ensembles are known to improve the predictive accuracy and analysis of future climate projections, consensus regarding how models are aggregated is lacking. In this study, we propose a probabilistic machine learning framework to systematically combine 13 regional climate models from the Coordinated Regional Downscaling Experiment (CORDEX) over High Mountain Asia. Our approach accounts for seasonal and spatial biases within the models, enabling the prediction of more faithful precipitation distributions. The framework is validated against gridded historical precipitation data and is used to generate projections for the near-future (2036–2065) and far-future (2066–2095) under RCP4.5 and RCP8.5 scenarios.
Keywords: High Mountain Asia, precipitation, ensemble learning, multi-model, regional climate models, probabilistic, mixture of experts, Gaussian process, CORDEX-WAS
@@ 1 Introduction
The glaciers and snowfields of High Mountain Asia (HMA), shown in Figure 1, store more frozen water than anywhere else in the world after the Arctic and Antarctic polar caps. These mountains provide fresh water for more than 1.9 billion people through Asia’s major rivers (Immerzeel et al., 2020). Despite the importance of these water resources, much remains unknown about their distribution and how they will evolve under climate change. The largest source of uncertainty for modelling future water security over HMA remains precipitation (Nie et al., 2021, Orr et al., 2022, Wester et al., 2023). Regional climate models (RCMs) are the main tool for understanding future precipitation over HMA (Maussion et al., 2014, Norris et al., 2017, 2019, Orr et al., 2017, Norris et al., 2020, Palazzi et al., 2015). However, complex orography and the lack of in situ hydrological observations, combined with limited RCM resolution and parametrisations, make precipitation difficult to model in this region (Tazi et al., 2024, 2025, Girona-Mata et al., 2024). To overcome these challenges and understand the range of possible changes in precipitation, different RCMs with a variety of physics and model parametrisations are run, and their outputs combined.
Such collections of climate models are known as multi-model ensembles. In general, multi-model ensembles improve the accuracy of historical predictions and provide a way to explore the uncertainties in model-based climate projections (Palmer et al., 2005, Stainforth et al., 2007, Kharin and Zwiers, 2002). However, the way in which the models are combined is controversial (Tebaldi and Knutti, 2007). This includes disagreements about how to measure the accuracy of the aggregation (Eyring et al., 2019), deal with ensemble outliers (Sanderson and Knutti, 2012), account for the proximity of model design(Masson and Knutti, 2011, Knutti, 2010), and deal with an infinite number of possible ensemble members (Tél et al., 2020). As a result, the Intergovernmental Panel on Climate Change has previously opted for the unweighted average of the global climate models (GCMs) from the Climate Model Intercomparison Project (CMIP) in their previous reports, using model spread as the uncertainty for future predictions (Flato et al., 2014). [^1] The sixth report (AR6) discards ‘hot-models’ from CMIP6 that have been shown to be overly sensitive to emission forcing (Zelinka et al., 2020, Hausfather et al., 2022, Lee et al., 2021).
@@
Second, we aggregate the RCM surrogate predictions at each spatiotemporal point via a mixture of experts (MoE), where distributions of experts (i.e. RCM surrogates) are combined using weights. We parametrise the weight of each expert by their Wasserstein distance with respect to historical data (APHRODITE) and a learnt ‘statistical temperature’. Our ensemble learning method is therefore similar to Bayesian model averaging (Massoud et al., 2023, Min et al., 2007, Olson et al., 2016), where the similarity of the probability distribution with respect to historical data is used as a measure of faithfulness. However, unlike previous ensembling techniques, our method neither gives each RCM one vote (i.e. equal weighting), nor does it discard outliers with less accurate distributions. Instead, we only down-weigh such outliers. As more experts are added, more precipitation events can be considered, and the final distribution refined.
The paper is structured as follows. The study area and data are first described in Section 2. The method is then summarised in Section 3 and validated in Section 4. Predictions for a historical reference period (1976-2005), the near future (2036-2065), and the far future (2066-2095) for RCP4.5 and RCP8.5 are analysed in Section 5. Finally, limitations and further work are discussed in Section 6.
Figure 1: Map of High Mountain Asia showing major rivers and lakes (Lehner and Grill, 2013), the spatial distribution of glaciers (RGI Consortium, 2017) and permafrost (Westermann et al., 2024), the HKH boundary as defined by ICIMOD (2008), and three mountain subregions used for this study; West Himalaya and Karakoram (HMA1), Central and East Himalaya (HMA2), the Hengduan Shan and Southeast Tibet (HMA3). The map also includes the standardised names of glacierised mountains regions (GTN-G, 2023) referenced in this paper.
@@ 2.0.2 CORDEX-WAS
CORDEX is a global initiative that aims to provide high-resolution climate data for regional and local applications. CORDEX for the West Asia domain (CORDEX-WAS, 20°-20°- 20  °-50°50°50  °N 20°20°20  °-115°115°115  °E), also known as the South Asia domain, uses RCMs to dynamically downscale global climate simulations from the fifth CMIP (CMIP5) over most of HMA (Sanjay et al., 2017, Taylor et al., 2012). CORDEX-WAS outputs are chosen over other RCM simulations for their accessibility and rigorous experimental protocol (Giorgi and Gutowski Jr, 2015), which allow for the straightforward comparison of many models. The CORDEX-WAS experiments are made up of 13 models using three RCMs and ten driving CMIP5 GCMs, as listed in Table 1.
Table 1: CORDEX-WAS RCM simulations with CMIP5 driving models. Note that GFDL-ESM2M, MPI-ESM-LR, and CNRM-CM5 each drive two RCMs. †The SMHI-RCA4 CNRM-CM5 model run only extends to run only extends to 2085 for RCP8.5. ‡MPI-REMO2009 MPI-ESM-LR model output only begins from 1961.
@@ 2.0.3 APHRODITE
The Asian Precipitation-Highly Resolved Observational Data Integration Towards Evaluation of Water Resources (APHRODITE) dataset is a gridded precipitation dataset ranging from 1951 to 2015 with a spatial resolution of 0.25°0.25°0.25  ° (Yatagai et al., 2012). In general, APHRODITE has one of the best spatiotemporal coverage of gridded precipitation products over High Mountain Asia and is one of the most studied and accurate products for the region (Dimri, 2021). The data set was created through the interpolation of precipitation gauge observations using a custom correlation distance lookup table. We use the gridded monthly precipitation product APHRO_V1101 specifically developed for monsoon Asia. Figure 2 plots APHRODITE precipitation between 1976 and 2005.
@@ 3 Method
Figure 3: Ensemble learning method. For a given emission scenario and climatological period, each of the $R$ RCMs outputs from CORDEX-WAS is split into $J$ manageable spatiotemporal domains, $\mathcal{D}^{(j)}_{r}$ for the $j$th domain and $r$th RCM. A GP is fit to each domain and then combined using BCM. The outputs of each RCM GP are then combined using a weighted mixture model or mixture-of-experts (MoE) with weights $\bm{w}_{r}$.
@@ 3.1.1 Gaussian processes
The problem is modelled as $y_{i}=f(\bm{x}_{i})+\epsilon_{i}$, where we assume that the noise $\epsilon_{i}\sim\mathcal{N}(\bm{0},\sigma_{\mathrm{n}}^{2})$ is independently and identically distributed. The function $f$ can be represented with a Gaussian process (GP) (Rasmussen and Williams, 2006). GPs are powerful non-parametric Bayesian models, often used for regression. A GP is defined as a distribution over functions such that every finite set of inputs and outputs is jointly Gaussian distributed. GPs are fully defined by a mean function $\mu(\cdot)$ and a kernel function $k(\cdot,\cdot)$, which depend on hyperparameters $\bm{\theta}$.
If we assume the function $f$ to be a GP, $f\sim\mathcal{GP}(\mu(\cdot),k(\cdot,\cdot))$, it follows that $f(\bm{X})\sim\mathcal{N}(\bm{\mu}_{\bm{X}},\bm{K}_{\bm{X}}+\sigma_{\mathrm{n}}^{2}\bm{I})$ where $(\bm{\mu}_{\bm{X}})_{i}:=\mu(\bm{x}_{i})$ is a $N$-dimensional vector and $(\bm{K}_{\bm{X}})_{ij}=k(\bm{x}_{i},\bm{x}_{j})$ is an $N\times N$ covariance matrix. We infer a posterior GP conditioned on the dataset $\mathcal{D}=(\bm{X},\bm{Y})$. The mean and variance of the Gaussian posterior predictive distribution of $f$, evaluated at a test point $\bm{x}_{*}$, are given by
@@
$\hat{\theta}=\mathop{\mathrm{argmax}}_{\theta}[\log p(\bm{Y}|\bm{X},\bm{\theta})]=\mathop{\mathrm{argmax}}_{\theta}\left[\log\mathcal{N}\left(\bm{Y}|\bm{\mu}_{\bm{X}},\bm{K}_{\bm{X}}+\sigma_{\mathrm{n}}^{2}\bm{I}\right)\right].$ (3)
Once a distribution is modelled as a GP, it can be evaluated at any inputs. Therefore, GPs can be used to interpolate empirical distributions. To compare the RCM surrogates with the APHRODITE data set Section 2.0.3, we evaluate the surrogate GPs at the APHRODITE grid points.
@@ 3.1.2 Warped GP
where precipitation $y>0$ is assumed to be positive and $\lambda$ is the scaling factor. The scaling factor is optimised to make the distribution of $g_{\lambda}(y)$ as Gaussian as possible by maximising the log-likelihood of the data originating from a Gaussian distribution (Box and Cox, 1964). In this study, we use a different optimised scaling factor $\lambda_{r}$ for each RCM $r$. Henceforth, we will use $f(\cdot)$ to denote the function that models the output warped $\tilde{y}$. We will denote the unwarped function, which may have heavy tails in its distribution, as $f_{\lambda}(\cdot):=g^{-1}_{\lambda}(f(\cdot))$. Details on the distribution of the unwarped precipitation can be found in B.
@@ 3.1.3 Chained GP
For this application, the variance of precipitation $y$ should be considered as the intrinsic climatological variance rather than the noise. The GP output should therefore be heterogeneous in its variance, i.e. heteroskedastic. To address these requirements, we use chained GP (Saul et al., 2016) to model the variance and mean as latent functions such that:
@@
where $f_{1}\sim\mathcal{GP}(0,k_{1}(\cdot,\cdot))$, $f_{2}\sim\mathcal{GP}(0,k_{1}(\cdot,\cdot))$ and $\alpha$ is a function used to map the outputs of $f_{2}$ to positive values. This mapping is required as the output of $\alpha$ represents the variance of a Gaussian distribution. Here, we apply an exponential function $\alpha(f_{2}(\bm{x}_{i}))=e^{f_{2}(\bm{x}_{i})}$, but other functions, such as softplus, could be used. For our work, use a shortcut to calculate the variance at each point without inducing points by exploiting the gridded nature of our data.
@@ 3.1.4 Bayesian committee machine
Due to GP’s poor scalability to large datasets, we use robust Bayesian committee machines (BCMs) as a way of combining multiple GPs and speeding up inference (Tresp, 2000, Deisenroth and Ng, 2015, Cohen et al., 2020). To do this, the training dataset is partitioned into $J$ subsets of size $M$ (so $N=JM$) where $M\ll N$. We define these subsets or domains as $\mathcal{D}^{(j)}=(\bm{X}^{(j)},\bm{Y}^{(j)})$ for $j=1,\dots,J$. The robust BCM GPs share kernel hyperparameters. This automatically regularises the population of experts: individual experts cannot overfit to the local subset of the data they are fed with due to this shared set of hyperparameters. Assuming conditional independence between domains $\mathcal{D}^{(j)}$ and by repeated application of Bayes’ theorem, the following predictive distribution is obtained:
@@
where $p(f(\bm{x}_{*}))$ is the prior distribution of the GP evaluated at test point $\bm{x}_{*}$ and $\beta_{j}$ controls the contribution of expert $j$ (see A). This distribution is inherently Gaussian as the product of Gaussian distributions is also Gaussian. The BCM’s predictive mean and precision (i.e. the inverse of the variance) are given by
@@
respectively. $\mu^{j}_{*}$ and $(\sigma^{j}_{*})^{-2}$ are mean and precision of the predictive posterior given domain $\mathcal{D}^{(j)}$, and $(\sigma^{\text{prior}}_{*})^{-2}$ is the prior precision of $p(f_{*})$. This setup conserves the properties of an exact GP while improving scalability, moving from $\mathcal{O}(N^{3})$ for training and $\mathcal{O}(N^{2})$ for prediction to $\mathcal{O}(JM^{3})$ and $\mathcal{O}(JM^{2})$ respectively (Deisenroth and Ng, 2015). For our work, the domains are divided temporally, such that each domain contains precipitation for a particular month over the whole spatial area. In Figure 4, we illustrate the typical surrogate output with CSIRO RegCM4 for February. The surrogate distributions closely match the empirical RCM histograms, which can significantly differ depending on time and location.

### v2

@@ Abstract
High Mountain Asia holds the largest concentration of frozen water outside the polar regions, serving as a crucial water source for more than 1.9 billion people. In the face of climate change, precipitation represents the largest source of uncertainty for hydrological modelling in this area. Future precipitation predictions remain challenging due to complex orography, lack of in situ hydrological observations, and limitations in climate model resolution and parametrisation for this region. To address the uncertainty posed by these challenges, climate models are often aggregated into multi-model ensembles. While multi-model ensembles are known to improve the predictive accuracy and analysis of future climate projections, consensus regarding how models are aggregated is lacking. In this study, we propose a probabilistic machine learning framework to combine 13 regional climate models from the Coordinated Regional Downscaling Experiment (CORDEX) over High Mountain Asia. Our approach accounts for seasonal and spatial biases within the models, enabling the prediction of more faithful precipitation distributions. The framework is validated against gridded historical precipitation data and is used to generate projections for the near future (2036–2065) and far future (2066–2095) under RCP4.5 and RCP8.5 scenarios.
Keywords: High Mountain Asia, precipitation, ensemble learning, regional climate models, probabilistic, mixture of experts, Gaussian process, CORDEX-WAS
@@ 1 Introduction
The glaciers and snowfields of High Mountain Asia (HMA), shown in Figure 1, store more frozen water than anywhere else in the world after the Arctic and Antarctic polar caps. These mountains provide fresh water for more than 1.9 billion people through Asia’s major rivers (Immerzeel et al., 2020). Despite the importance of these water resources, much remains unknown about their distribution and how they will evolve under climate change. Precipitation remains he largest source of uncertainty for modelling future water security over HMA (Nie et al., 2021, Orr et al., 2022, Wester et al., 2023). Regional climate models (RCMs) are the main tool for understanding future precipitation over HMA (Maussion et al., 2014, Norris et al., 2017, 2019, Orr et al., 2017, Norris et al., 2020, Palazzi et al., 2015). However, complex orography and the lack of in situ hydrological observations, combined with limited RCM resolution and parametrisations, make precipitation difficult to model in this region (Tazi et al., 2024, 2025, Girona-Mata et al., 2024). To overcome these challenges and understand the range of possible changes in precipitation, different RCMs with different physics and model parametrisations are run, and their outputs combined.
Such collections of climate models are known as multi-model ensembles. In general, multi-model ensembles improve the accuracy of historical predictions and provide a way to explore the uncertainties in model-based climate projections (Palmer et al., 2005, Stainforth et al., 2007, Kharin and Zwiers, 2002). However, the way in which models are combined is controversial (Tebaldi and Knutti, 2007). This includes disagreements about how to measure the aggregate accuracy (Eyring et al., 2019), deal with ensemble outliers (Sanderson and Knutti, 2012), account for the proximity of model design (Masson and Knutti, 2011, Knutti, 2010), and address an infinite number of possible ensemble members (Tél et al., 2020). As a result, the Intergovernmental Panel on Climate Change has previously opted for the unweighted average of the global climate models (GCMs) from the Climate Model Intercomparison Project (CMIP) in their previous reports, using model spread as the uncertainty for future predictions (Flato et al., 2014). [^1] The sixth report (AR6) discards ‘hot-models’ from CMIP6 that have been shown to be overly sensitive to emission forcing (Zelinka et al., 2020, Hausfather et al., 2022, Lee et al., 2021).
@@
Second, we aggregate the RCM surrogate predictions at each spatiotemporal point via a mixture of experts (MoE), where distributions of experts (i.e. RCM surrogates) are combined using weights. We parametrise the weight of each expert by their Wasserstein distance with respect to historical data (APHRODITE) and a learnt ‘statistical temperature’. Our ensemble learning method is therefore similar to Bayesian model averaging (Massoud et al., 2023, Min et al., 2007, Olson et al., 2016), where the similarity of the probability distribution with respect to historical data is used as a measure of faithfulness. However, unlike previous ensembling techniques, our method neither gives each RCM one vote (i.e. equal weighting), nor does it completely discard outliers with less accurate distributions, only down-weigh them. As more experts are added, more precipitation events can be considered, and the final distribution refined.
The paper is structured as follows. The study area and data are first described in Section 2. The method is then summarised in Section 3 and validated in Section 4. Predictions for a historical reference period (1976–2005), the near future (2036–2065), and the far future (2066–2095) for RCP4.5 and RCP8.5 are analysed in Section 5. Finally, limitations and further work are discussed in Section 6.
Figure 1: Map of High Mountain Asia showing major rivers and lakes (Lehner and Grill, 2013), glaciers (RGI Consortium, 2017), permafrost (Westermann et al., 2024), the HKH boundary as defined by ICIMOD (2008), and three mountain subregions used for this study: the West Himalaya and Karakoram (HMA1), the Central and East Himalaya (HMA2), the Hengduan Shan and Southeast Tibet (HMA3). The map also includes the standardised names of glacierised mountains regions (GTN-G, 2023) referenced in this paper.
@@ 2.0.2 CORDEX-WAS
CORDEX is a global initiative that aims to provide high-resolution climate data for regional and local applications. CORDEX for the West Asia domain (CORDEX-WAS, 20°-20°- 20  °50°50°50  °N 20°20°20  °115°115°115  °E), also known as the South Asia domain, uses RCMs to dynamically downscale global climate simulations from the fifth CMIP (CMIP5) over most of HMA (Sanjay et al., 2017, Taylor et al., 2012). CORDEX-WAS outputs are chosen over other RCM simulations for their accessibility and rigorous experimental protocol (Giorgi and Gutowski Jr, 2015), which allow for the straightforward comparison of many models. CORDEX-WAS is made up of 13 models using three RCMs and ten driving CMIP5 GCMs, as listed in Table 1.
Table 1: CORDEX-WAS RCMs with CMIP5 driving models. Note that GFDL-ESM2M, MPI-ESM-LR, and CNRM-CM5 each drive two RCMs. †The SMHI-RCA4 CNRM-CM5 model output only extends to 2085 for RCP8.5. ‡MPI-REMO2009 MPI-ESM-LR model output only begins from 1961.
@@ 2.0.3 APHRODITE
The Asian Precipitation-Highly Resolved Observational Data Integration Towards Evaluation of Water Resources (APHRODITE) dataset is a gridded precipitation dataset ranging from 1951 to 2015 with a spatial resolution of 0.25°0.25°0.25  ° (Yatagai et al., 2012). In general, APHRODITE has one of the best spatiotemporal coverage of precipitation products over High Mountain Asia and is one of the most studied and accurate products for the region (Dimri, 2021). The dataset was created through the interpolation of precipitation gauge observations using a custom correlation distance lookup table. We use the monthly precipitation product APHRO_V1101 specifically developed for monsoon Asia. Figure 2 plots APHRODITE precipitation between 1976 and 2005.
@@ 3 Method
Figure 3: Ensemble learning method. For a given emission scenario and climatological period, each of the $R$ RCMs outputs from CORDEX-WAS are split into $J$ manageable spatiotemporal domains, $\mathcal{D}^{(j)}_{r}$ for the $j$th domain and $r$th RCM. A GP is fit to each domain and then combined using a BCM. The outputs of each RCM GP are then combined using a weighted mixture model or mixture-of-experts (MoE) with weights $\bm{w}_{r}$.
@@ 3.1.1 Gaussian processes
The problem is modelled as $y_{i}=f(\bm{x}_{i})+\epsilon_{i}$, where we assume that the noise $\epsilon_{i}\sim\mathcal{N}(\bm{0},\sigma_{\mathrm{n}}^{2})$ is independently and identically distributed. The function $f$ can be represented with a Gaussian process (GP) (Rasmussen and Williams, 2006). GPs are powerful non-parametric Bayesian models. A GP is defined as a distribution over functions such that every finite set of inputs and outputs is jointly Gaussian distributed. GPs are fully defined by a mean function $\mu(\cdot)$ and a kernel function $k(\cdot,\cdot)$, which depend on hyperparameters $\theta_{\mu}$ and $\theta_{k}$, jointly referred to as $\theta$.
If we assume $f$ to be a GP, i.e. $f\sim\mathcal{GP}(\mu(\cdot),k(\cdot,\cdot))$, it follows that $f(\bm{X})\sim\mathcal{N}(\bm{\mu}_{\bm{X}},\bm{K}_{\bm{X}}+\sigma_{\mathrm{n}}^{2}\bm{I})$ where $(\bm{\mu}_{\bm{X}})_{i}:=\mu(\bm{x}_{i})$ is a $N$-dimensional vector and $(\bm{K}_{\bm{X}})_{ij}=k(\bm{x}_{i},\bm{x}_{j})$ is an $N\times N$ covariance matrix. We infer a posterior GP conditioned on the dataset $\mathcal{D}=(\bm{X},\bm{Y})$. The mean and variance of the posterior predictive distribution of $f$, evaluated at a test point $\bm{x}_{*}$, are given by
@@
$\hat{\theta}=\mathop{\mathrm{argmax}}_{\theta}[\log p(\bm{Y}|\bm{X},\theta)]=\mathop{\mathrm{argmax}}_{\theta}\left[\log\mathcal{N}\left(\bm{Y}|\bm{\mu}_{\bm{X}}(\theta_{\mu}),\bm{K}_{\bm{X}}(\theta_{K})+\sigma_{\mathrm{n}}^{2}\bm{I}\right)\right].$ (3)
Once a distribution is modelled as a GP, it can be evaluated at any inputs. Therefore, GPs can be used to interpolate empirical distributions. To compare the RCM surrogates with the APHRODITE dataset introduced in Section 2.0.3, we evaluate the surrogate GPs at the APHRODITE grid points.
@@ 3.1.2 Warped GP
where precipitation $y>0$ is assumed to be positive and $\lambda$ is the scaling factor. The scaling factor is optimised to make the distribution of $g_{\lambda}(y)$ as Gaussian as possible by maximising the log-likelihood of the model parameters arising from placing a Gaussian distribution over the transformed observations (Box and Cox, 1964). In this study, we use a different optimised scaling factor $\lambda_{r}$ for each RCM $r$. Henceforth, we will use $f(\cdot)$ to denote the function that models the warped output $\tilde{y}$. We will denote the unwarped function, which may have heavy tails in its distribution, as $f_{\lambda}(\cdot):=g^{-1}_{\lambda}(f(\cdot))$. Details on the distribution of the unwarped precipitation can be found in B.
@@ 3.1.3 Chained GP
For this application, the variance of precipitation $y$ should be considered as intrinsic climatological variance rather than noise. The GP output should therefore be heterogeneous in its variance, i.e. heteroskedastic. To address these requirements, we use chained GPs (Saul et al., 2016) to model the variance and mean as latent functions such that
@@
where $f_{1}\sim\mathcal{GP}(0,k_{1}(\cdot,\cdot))$, $f_{2}\sim\mathcal{GP}(0,k_{1}(\cdot,\cdot))$ and $\alpha$ is a function used to map the outputs of $f_{2}$ to positive values. This mapping is required as the output of $\alpha$ represents the variance of a Gaussian distribution. Here, we apply an exponential function $\alpha(f_{2}(\bm{x}_{i}))=e^{f_{2}(\bm{x}_{i})}$, but other functions, such as softplus, could be used. For our work, we use a shortcut to calculate the variance at each point without using further approximations, such as inducing points (Titsias, 2009), by exploiting the gridded nature of our data.
@@ 3.1.4 Bayesian committee machine
Due to GP’s poor scalability to large datasets, we use robust Bayesian committee machines (BCMs) as a way of combining multiple GPs and speeding up inference (Tresp, 2000, Deisenroth and Ng, 2015, Cohen et al., 2020). To do this, the training dataset is partitioned into $J$ subsets of size $M$ (so $N=JM$) where $M\ll N$. We define these subsets or domains as $\mathcal{D}^{(j)}=(\bm{X}^{(j)},\bm{Y}^{(j)})$ for $j=1,\dots,J$. GPs are then used to model each domain. The robust BCM GPs share kernel hyperparameters, stopping individual experts from overfitting their local subset of data. Assuming conditional independence between domains $\mathcal{D}^{(j)}$ and by repeated application of Bayes’ theorem, the following predictive distribution is obtained by combining the separate GP models:
@@
where $p(f(\bm{x}_{*}))$ is the prior distribution of the GP evaluated at test point $\bm{x}_{*}$ and $\beta_{j}$ controls the contribution of expert $j$ (see A). The robust BCM’s predictive mean and precision (inverse of the variance) are given by
@@
respectively, where $\mu^{j}_{*}$ and $(\sigma^{j}_{*})^{-2}$ are the mean and precision of the predictive posterior given domain $\mathcal{D}^{(j)}$, and $(\sigma^{\text{prior}}_{*})^{-2}$ is the prior precision of $p(f_{*})$. This setup conserves the properties of an exact GP while improving scalability, moving from $\mathcal{O}(N^{3})$ for training and $\mathcal{O}(N^{2})$ for prediction to $\mathcal{O}(JM^{3})$ and $\mathcal{O}(JM^{2})$ respectively (Deisenroth and Ng, 2015). For our work, the domains are divided temporally, such that each domain contains precipitation for a particular month over the whole spatial area. In Figure 4, we illustrate the typical surrogate output with CSIRO RegCM4 for February. The surrogate distributions closely match the empirical RCM histograms , which can significantly differ depending on time and location.